3. (a) A coach buys 3 cricket bats and 6 balls for
3,900. Later, he buys 1 bat and 3 balls for
another team for 1,450. Write two linear
equations to represent the purchases. Why
can the linear system so obtained be solved
by Cramer's rule ? Also, use the rule to
solve the system, and interpret the solution
in the given context. 4
(b) Give an example of a 4 x 1 matrix with
entries from Z \ N.
1
Expert's answer
2019-11-18T10:32:53-0500

3 (a). 3x+6y=3900

x+3y=1450



"\Delta=\begin{vmatrix}
 3 & 6 \\
 1 & 3
\end{vmatrix}=9-6=3"

is not equal to zero;

"\Delta x=\begin{vmatrix}
 3900& 6 \\
 1450 & 3
\end{vmatrix}=3900*3-1450*6=3000"

"\Delta y=\begin{vmatrix}
 3 & 3900 \\
 1 & 1450
\end{vmatrix}=3*1450-1*3900=450"

"x=\Delta x/\Delta=3000/3=1000"

"y=\Delta y/ \Delta=450/3=150"

x stands for bats, y stands for balls.


4 (b).

"A=\begin{bmatrix}
 -1 \\
 -2 \\
-3 \\
-4
\end{bmatrix}"